The mean distance of Mars from the Earth is 0.5 A.U. and its orbital period is 687 days. Calculate the orbital period of Jupiter given that its mean distance from the Earth is 4 A.U.
1
Expert's answer
2019-04-01T10:40:44-0400

According to the Third Kepler’s Law, the orbital period of a planet is directly proportional to the cube of the semi-major axis of its orbit. In this case, we can write


"\frac{T_M^2}{a_M^3} = \frac{T_J^2}{a_J^3} (1)"

The axis of Mars aM is 1.5 A. U. (0.5 A. U. form the Earth plus 1 A. U. between the Earth and the Sun) while the axis of Jupiter aJ is 5 A. U. (4 A. U. form the Earth plus 1 A. U. between the Earth and the Sun).


Using (1), we got:


"T_J= \sqrt{\frac{T_M^2\times a_J^3}{a_M^3}} (2)"

We calcuted using (2): TJ=4181 days


Answer:

4181 days